5x + 6y=28
Find x & y
I need a solution to this problem.. How to solve this problem.
1
Expert's answer
2012-10-19T11:46:41-0400
we can rewrite x=1/5(28-6y)=5-y+(3-y)/5
so we can get many solutions for
every y we get some x
if we want to get integer solutions then y=3+5k,
x=5-3-5k-k=2-6k for integer k